OpenStudy (anonymous):

Part 1: Change 2x + 5y = –30 to slope-intercept form. You must show all work. Part 2: State the slope and y-intercept

OpenStudy (anonymous):

How do i change this to slope intercept form ?

OpenStudy (yrelhan4):

Write the equation in the form of y=mx+c.. Thats the slope intercept form.

OpenStudy (anonymous):

i thought it was y= mx+b what is c ?

OpenStudy (yrelhan4):

They are the same. b or c doesnt matter.

OpenStudy (anonymous):

can u put it in the form for me and i will try to solve . pls

OpenStudy (yrelhan4):

2x + 5y = –30 5y = -30 -2x y = -6 - (2/5)x or y = -2/5x - 6

OpenStudy (anonymous):

did u just solve it lol ?

OpenStudy (yrelhan4):

Eh, i just wrote in the form of y = mx + b..